OpenStudy (anonymous):

find the values of c guaranteed by the Mean Value Theorem for Integrals for the function over the given interval f(x)=x^3 i am lost, i am trying to take the integral, then solving it but i have i gotten the wrong answer right , i have been using the mean vaule theorem but it is not qorking please help

OpenStudy (anonymous):

on what interval?

OpenStudy (anonymous):

[0,3]

OpenStudy (anonymous):

what does the theorem say?

OpenStudy (anonymous):

it says there exist a number c in [0,3] such that \[f(c)(b-a) = \int\limits_{a}^{b} f(x) dx\]

OpenStudy (anonymous):

a = 0 b = 3 f(x) = x^3 solve for c

OpenStudy (anonymous):

if f is integrable [a,b] then the average value is the function on [a,b] os avg. 1/b-a \[\int\limits_{a}^{b}f(x)dx\]

OpenStudy (anonymous):

the average value is f(c). In other words, it's the y value, not x. c is an x value in [0,3]

OpenStudy (anonymous):

you're supposed to find c. Not f(c)

OpenStudy (anonymous):

i have been trying to use the theorem, like i see in my notes i get something else then the answer in the back which is 3\[3\sqrt[3]{2}/2\]

OpenStudy (anonymous):

answer should be 2 (2/3)^(1/3)

OpenStudy (anonymous):

you probably looked at the solution to a different problem. Either that or, you copied the wrong question

OpenStudy (anonymous):

no its the same quesition and answer

OpenStudy (anonymous):

maybe i looks at the wrong formula hang on \[\int\limits_{a}^{b}f(x)dx=f(c)-[b-a]\]

OpenStudy (anonymous):

i made a mistake. But the correct answer is still different from yours https://www.wolframalpha.com/input/?i=c%5E3+%283-0%29+%3D+%281%2F4%29+%283%29%5E4

OpenStudy (anonymous):

yeah, your formula is wrong

OpenStudy (anonymous):

it's f(c)(b-a). not f(c) - (b-a)

OpenStudy (anonymous):

ok thats where i am going wrong thank you